OpenStudy (anonymous):

A population is growing at a rate of 2 ,4, 8, 16, 32. Which type of growth does this describe? A. linear B. logistic C. exponential D. consistent

OpenStudy (aaronq):

Can you identify the pattern of growth?

OpenStudy (aaronq):

2 2*2=4 2*2*2=8 2*2*2*2=16 2*2*2*2*2=32 \(\sf \large 2^n\)